OpenStudy (anonymous):

what are the zeroes of the fuction? x(squared)-2x+5=0 Fan and medal!

OpenStudy (freckles):

you could use the quadratic formula

OpenStudy (anonymous):

ok

OpenStudy (freckles):

\[ax^2+bx+c=0 \\ x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] steps: identify and plug in and evaluate

OpenStudy (anonymous):

\[x=\div2\pm \sqrt{-2^2-4(1)(5)}\]

OpenStudy (anonymous):

over 2(1)

OpenStudy (anonymous):

@freckles I got \[x=2\pm \sqrt{16}\div2\]

OpenStudy (freckles):

\[x=\frac{-(-2) \pm \sqrt{(-2)^2-4(1)(5)}}{2(1)}=\frac{2 \pm \sqrt{4-20}}{2}=\frac{2 \pm \sqrt{-16}}{2}\] close

OpenStudy (freckles):

this can be simplified though